OpenStudy (anonymous):

What is the total amount due for a loan of $824 at an interest rate of 7.5% for 9mths. I need the formula using I=P R T. can someone help me please.

OpenStudy (anonymous):

given p= 824, r= 7.5% t = 9/12 I= 824 x 7.5/100 x 9/12 I= 46.35

OpenStudy (anonymous):

Is it simple interest or compound interest??

OpenStudy (anonymous):

i believe simple

OpenStudy (anonymous):

Ok, then reckkimini's ans is correct, u can use it

OpenStudy (anonymous):

well think you for double checking I really do appreciate it.

OpenStudy (anonymous):

welcome, all credit to reckkimini...